### v3

# Empowering Agentic Video Analytics Systems with Video Language Models

Abstract.
AI-driven video analytics has become increasingly important across diverse domains. However, existing systems are often constrained to specific, predefined tasks, limiting their adaptability in open-ended analytical scenarios. The recent emergence of Video-Language Models (VLMs) as transformative technologies offers significant potential for enabling open-ended video understanding, reasoning, and analytics. Nevertheless, their limited context windows present challenges when processing ultra-long video content, which is prevalent in real-world applications. To address this, we introduce Avas, a VLM-powered system designed for open-ended, advanced video analytics. Avas incorporates two key innovations: (1) the near real-time construction of Event Knowledge Graphs (EKGs) for efficient indexing of long or continuous video streams, and (2) an agentic retrieval-generation mechanism that leverages EKGs to handle complex and diverse queries. Comprehensive evaluations on public benchmarks, LVBench and VideoMME-Long, demonstrate that Avas achieves state-of-the-art performance, attaining 62.3% and 64.1% accuracy, respectively, significantly surpassing existing VLM and video Retrieval-Augmented Generation (RAG) systems. Furthermore, to evaluate video analytics in ultra-long and open-world video scenarios, we introduce a new benchmark, Avas-100. This benchmark comprises 8 videos, each exceeding 10 hours in duration, along with 120 manually annotated, diverse, and complex question-answer pairs. On Avas-100, Avas achieves top-tier performance with an accuracy of 75.8%.
The Work is done during internship at Microsoft Research
copyright: acmcopyrightcopyright: rightsretainedconference: Preprint; price: ;
1. Introduction
Video analytics (Jiang et al., 2021a; Bhardwaj et al., 2022; Khani et al., 2023; Nan et al., 2023) has emerged as a transformative technology across a wide array of domains, such as security surveillance, intelligent transportation systems, industrial automation, and retail monitoring. By harnessing the power of deep learning models, video analytics systems extract meaningful insights, identify patterns, and generate actionable information from video data, enabling organizations to monitor, detect, and respond to events with greater efficiency and precision. The desired features of video analytics systems necessitate a variety of capabilities. We classify the intelligence levels of existing and future video analytics systems into five levels, labeled as L1 through L5:
L1 systems for specific classification, segmentation, and detection using models e.g., ResNet (He et al., 2016) and EfficientDet (Tan et al., 2020) to extract spatial information from video data, including object classes and bounding boxes, etc (Jiang et al., 2021a; Bhardwaj et al., 2022; Khani et al., 2023).
L2 systems go beyond spatial information extraction by enabling causal event detection and analytics, i.e., identifying short-term events. They use models like C3D (Tran et al., 2015) and ActionFormer (Zhang et al., 2022) to detect and localize events (e.g., actions, activities, anomalies) through spatiotemporal modeling.
L3 systems extend L2’s spatiotemporal detection by incorporating neural language processing (NLP). Using models like CLIPBERT (Lei et al., 2021), they interpret and respond to natural language queries (e.g., ”What animals appeared in the videos?”) instead of merely outputting fixed labels. Though queries are still limited to specific domains, L3 systems greatly improve user interaction for video analytics.
Despite significant advancements, current video analytics systems (Khani et al., 2023; Bhardwaj et al., 2022; Zhang et al., 2024b; Padmanabhan et al., 2023; Lu et al., 2022; Sivaraman et al., 2024; Zhang et al., 2021; Jiang et al., 2021a) primarily focus on L1 to L3 levels. These systems are designed for closed-end analytics, relying on specialized models for specific tasks, which limits their flexibility and adaptability. Consequently, we envision L4 and L5 systems to enable open-end analytics.
L4 systems enable open-ended video comprehension, reasoning, and analytics, marking a major advancement. They support general-purpose analytics on video sources, handling natural language queries and responses while enabling complex, long-term spatiotemporal reasoning. For example, they can answer questions like: ”What abnormal events occurred in the past ten hours?”, ”What caused the person to fall?”, or ”How did the animals behave after appearing on camera?”
L5 systems advance L4 by autonomously integrating external public or domain-specific knowledge to uncover explicit and implicit connections between videos and world knowledge, fulfilling the ultimate goal of video analytics: deriving meaningful insights and offering actionable automations.
In this paper, we delve into the development of L4 video analytics systems, leveraging the transformative potential of video-language models (VLMs). By combining vision and language understanding, VLMs enable generalized visual detection and advanced video comprehension, including causal reasoning, key information retrieval, and human-interpretable explanations. The integration of perception, reasoning, and interaction makes VLMs highly adaptable, positioning them as a key foundation for L4 systems.
However, integrating VLMs into video analytics poses significant challenges, primarily due to the limited context window of current VLMs compared to the extensive duration of video sources in typical video analytics scenarios. While L1 to L3 systems handle spatial and short-term causal event detection by processing frames independently or within small sliding windows (e.g., a few seconds), L4 systems require collective analysis of related frames for long-term causal detection, summarization, and reasoning. Current VLMs, like QwenVL (Wang et al., 2024a), can processing up to 768 frames, covering minutes or hours of video. However, video analytics often involves much longer sources, spanning hundreds of hours or continuous streams, far beyond the capabilities of existing VLMs.
Recent studies (Shen et al., 2024, 2025; Wang et al., 2025) that attempt to extend the context window of VLMs remain inadequate for processing video sources spanning hundreds of hours. Retrieval-augmented generation (RAG) frameworks (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025; Research, 2024) aim to address similar limitations by first retrieving relevant frames from massive contents and then generating final answers. However, these approaches still face significant challenges in handling the video modality (as detailed in §7), leading to both reduced analytics accuracy and substantial computational overhead.
In this paper, we present Avas, a system that integrates VLMs into video analytics to enable L4 capabilities. The core innovation of Avas is its efficient indexing mechanism, designed to handle extremely long video sources or unlimited video streams, and by leveraging the index, Avas effectively retrieves relevant information for a given query and generates accurate, robust responses. Specifically, Avas introduces the following key features that distinguish it from existing systems: 1) analyzing extremely long videos, spanning hundreds of hours or even unlimited video streams; 2) supporting for near-real-time (e.g., at more than 1 FPS) index construction and analytics; and 3) handling diverse and complex queries including temporal grounding and reasoning, summarization, event and entity understanding, and key information retrieval.
Particularly, Avas introduces two key components: near-real-time index construction and agentic retrieval and generation. During the index construction phase (§4), we propose event knowledge graphs (EKGs) as an indexing mechanism for video analytics. Unlike traditional knowledge graphs (KGs) used in text-based RAG systems (Edge et al., 2024; Guo et al., 2024; Research, 2024), EKGs represent a flow of insightful events, effectively capturing video dynamics and temporal consistency. Using a small VLM, such as QwenVL-7B, Avas extracts information to construct the EKG. To optimize this process, we introduce techniques that enable near-real-time index construction, achieving more than 5 FPS on typical edge servers equipped with 2 $\times$ 4090 GPUs.
In the retrieval and generation phase (§5), Avas employs an agentic search mechanism instead of directly retrieving information from the constructed index. This approach allows Avas to proactively retrieve more relevant information by utilizing contextual hints captured within the EKG, enables Avas to handle complex queries, including summarization and multi-hop reasoning etc. Furthermore, we introduce techniques to enhance Avas to robustly generate answers based on the retrieved information.
We evaluate Avas on multiple public video understanding benchmarks, including LVBench (Wang et al., 2024b) and VideoMME-Long (Fu et al., 2024). These benchmarks collectively comprise approximately 400 videos and 2,500 questions. We compare Avas against a diverse range of baselines, including state-of-the-art (SOTA) VLMs such as GPT-4o (Achiam et al., 2023), Gemini (Team et al., 2023), Phi-4-Multimodal (Abdin et al., 2024a), Qwen2.5-VL-7B (Wang et al., 2024a), InternVL2.5-8B (Chen et al., 2024), and LLaVA-Video-7B (Zhang et al., 2024a), as well as typical video-RAG systems such as VideoTree (Wang et al., 2024c), VideoAgent (Wang et al., 2024d), DrVideo (Ma et al., 2024), and VCA (Yang et al., 2024). On both benchmarks, Avas establishes new SOTA performance, achieving 62.3% on LVBench and 64.1% on VideoMME-Long, respectively, significantly surpassing baselines by up to 16.9% and 5.2%.
In addition to the public video understanding benchmarks, we introduce a new benchmark, Avas-100, specifically designed to evaluate L4 video analytics tasks. Avas-100 comprises 8 ultra-long videos, each exceeding 10 hours in duration, and includes a total of 120 manually annotated questions and answers. The videos are carefully curated from typical video analytics scenarios, and the questions cover multiple key categories relevant to video analytics systems. Evaluation results show that Avas achieves significantly better performance on Avas-100 compared to various baselines, with improvements of approximately 20.8%. In summary, we make the following contributions in this paper:
  • We propose Avas, the first L4 video analytics system powered by VLMs, to the best of our knowledge.
  • Avas introduces near-real-time index construction and agentic retrieval and generation, along with innovative techniques that enable key features for L4 video analytics, including open-ended analytics on extremely long video sources in near-real-time.
  • We evaluate Avas on two widely-used video understanding benchmarks, achieving SOTA performance with 62.3% on LVBench and 64.1% on VideoMME-Long.
  • Furthermore, we present Avas-100, a benchmark specifically designed for L4 video analytics systems, where Avas demonstrates significant improvements, outperforming baselines by approximately 20.8%.
2. Related Work and Motivation
2.1. Video Analytics System and VLMs
The field of video analytics (VA) has seen significant advancements in recent years (Khani et al., 2023; Bhardwaj et al., 2022; Zhang et al., 2024b; Padmanabhan et al., 2023; Lu et al., 2022; Sivaraman et al., 2024; Zhang et al., 2021). Leveraging emerging deep learning models, modern VA systems can extract insightful information, such as object locations or counts, from video streams processed on edge or cloud servers.
Existing VA systems predominantly support closed-end analytics (L1 to L3 systems as mentioed in §1), often relying on shallow neural networks to extract predefined, task-specific, and constrained information. For instance, Remix (Jiang et al., 2021a) leverages fine-tuned EfficientDet (Jiang et al., 2021b) to generate bounding boxes for pedestrians. Consequently, the flexibility and adaptability of VA systems are fundamentally limited by the inherent constraints of the specific models they employ.
Recently, video language models (VLMs), e.g., GPT-4o (Achiam et al., 2023), Gemini (Team et al., 2023, 2024), QwenVL (Wang et al., 2024a; Bai et al., 2025) and Phi (Abdin et al., 2024a, b; Abouelenin et al., 2025), have demonstrated their transformative potential in video analytics tasks. By leveraging the extensive world knowledge embedded in large language models (LLMs), VLMs not only achieve generalized visual grounding but also exhibit advanced video comprehension capabilities, such as zero-shot temporal and spatial reasoning, contextual retrieval, and semantic understanding. More importantly, VLMs enable natural language interaction, allowing users to dynamically query video content. This makes them particularly well-suited for addressing diverse and unstructured open-ended analytics in real-world scenarios, towards L4 VA systems.
However, adopting VLMs in VA systems is far from straightforward. Existing L1 to L3 VA systems, which rely on traditional DNNs, typically process each video frame independently. In contrast, VLMs require related frames to be processed collectively to infer causal relationships and temporal dependencies across frames. This shift introduces significant complexity, as existing VLMs are generally capable of handling only minute-level or sub-hour-level videos due to the limited context window inherent in language models.
In real-world video analytics scenarios, the scale of videos to be analyzed is often vastly larger—spanning hundreds of hours or more (e.g., monitoring wildlife behavior over an entire month, as illustrated in Fig. 2). This creates a fundamental gap between the capabilities of VLMs and the demands of VA systems, as the limited context window length of VLMs directly restricts their ability to process videos of such extensive durations effectively.
2.2. Long Video Understandings
Recent research also increasingly focuses on enabling long video understandings (Shen et al., 2024; Wang et al., 2025). Given the inherent limitations of autoregressive language models, such as the constrained context window length, efforts have been directed toward reducing the redundancy in video inputs to facilitate the processing of extended video durations. For instance, LongVU (Shen et al., 2024) and AdaRETAKE (Wang et al., 2025) introduce dynamic compression mechanisms that prioritize video content based on its relevance, selectively retaining frames or regions most pertinent to downstream language tasks. Similarly, NVILA (Liu et al., 2024) addresses the efficiency-accuracy trade-off by optimizing sampling strategies and resolution to fit within limited token budgets.
While these approaches have succeeded in increasing the number of frames that models can process and mitigating the constraints of context windows to some extent, they fall short of achieving a fundamental breakthrough. Most existing methods remain limited, with the supported video length typically capped at approximately one hour (Qian et al., 2024), which is insufficient to meet the demands of video analytics. Moreover, as video length extends, the inference cost rises correspondingly, further exacerbating the challenges of scaling these systems.
2.3. Retrieval Augmented Generation
Figure 1. The system overview of Avas.
Short (1.4 minutes) Medium (9.7 minutes) Long (39.7 mintues)
Total Needed Total Needed Total Needed
2144.8 12.1 (0.5%) 13924.1 68.1 (0.4%) 66847.1 82.3 (0.1%)
Table 1. A small portion of frames is required to answer each specific question across the short, medium, and long video subsets of the VideoMME (Fu et al., 2024) benchmark using Qwen2-VL.
While the videos to be analyzed may span extensive durations, the frames necessary to respond to a specific query are often limited. To validate this observation, we conducted an experiment on VideoMME (Ma et al., 2024) using Qwen2-VL (Wang et al., 2024a). Specifically, we first identified all questions for which Qwen2-VL produced correct answers by uniformly sampling frames from the videos at a rate of 1 FPS. For these questions, we then determined the minimal set of frames required for the VLM to generate the correct answer by iteratively reducing the number of input frames using a binary search strategy  [^1] For example, we initially uniformly sample 100 frames from the video. If the VLM can generate the correct answer based on these frames, we then attempt to reduce the frame set to 50 frames. If 50 frames are still sufficient to produce the correct answer, we further reduce the set to 25 frames. Conversely, if 50 frames are insufficient, we increase the set to 75 frames, iteratively refining the frame selection using binary search strategy.. The results reveal that the frames necessary to answer a given question constitute only a small fraction of the total frames in the video, as shown in Table 1.
Based on this observation, an intuitive approach would be to first retrieve the relevant frames corresponding to a specific query and then generate the final answer based on these frames, a method commonly referred to as Retrieval-Augmented Generation (RAG). To retrieve potentially relevant frames, a straightforward strategy involves vectorized retrieval, where each frame of the video is embedded using a vision-language model, such as CLIP (Radford et al., 2021). At query time, the embedding of the query is used to retrieve relevant frames by comparing the similarity between the query embedding and the vectorized frames.
However, the vectorized retrieval method’s limitations stem from the detailed information contained in the query. Notably, such an approach struggles to handle query-focused summaries (Edge et al., 2024) (e.g., ”What happened in the last few hours?”) or multi-hop queries (Zhao et al., 2024) (e.g., ”What did the man do after he opened the fridge?”), as the retrieved frames often fail to capture key contexts that are not explicitly mentioned in the query descriptions.
To enable effective retrieval, recent research has explored two prominent approaches: video structuring and iterative retrieval. For example, Video-RAG (Luo et al., 2024) structures videos by utilizing various tools to extract information such as audio transcripts (ASR), optical character recognition (OCR) results, and object detection outputs. It then applies RAG techniques to the structured information. However, this method is inherently constrained by the tools employed for video structuring. It is often impractical to predict in advance what types of information need to be extracted and what corresponding tools should be utilized, limiting its adaptability to diverse and dynamic video analytics scenarios.
Alternatively, researchers have proposed obtaining relevant frames through multiple iterative retrieval processes (Wang et al., 2024d, c; Ma et al., 2024; Yang et al., 2024). For instance, VideoAgent (Wang et al., 2024d) typically begins with a coarse-grained sampling of video segments to establish an initial high-level understanding. Based on this, the VLM is prompted to decide which finer-grained segments to retrieve and analyze in subsequent iterations. However, these approaches face significant challenges when applied to video analytics scenarios involving extremely long videos. On one hand, the initial coarse-grained sampling may become insufficient as video length increases, potentially missing critical information. On the other hand, the iterative retrieval and analysis process becomes increasingly computationally expensive as video duration grows, making it impractical for large-scale video analytics tasks.
Recent studies have advanced RAG techniques (Edge et al., 2024; Research, 2024; Guo et al., 2024; Fan et al., 2025) by incorporating knowledge graph construction to enhance the retrieval process. However, these works primarily focus on text-only RAG problems, and adapting such approaches to video analytics remains a significant challenge due to the complexity and multimodal nature of video data. In this paper, we propose Avas, which, to the best of our knowledge, is the first system to enable VLM-powered video analytics by effectively addressing the aforementioned challenges.
3. Avas System Overview
The key idea of Avas lies in leveraging a small VLM to efficiently structure video streams into discrete events, then linking these events by extracting insights from each to construct a comprehensive index. Given a specific query, Avas leverages this index to proactively retrieve relevant information from both the index and the associated raw frames. Ultimately, the retrieved data are utilized by the VLM to produce a coherent and contextually appropriate response.
To build Avas as the L4 video analytics system, we established the following design principles: 1) The analytics should be scalable to any volume of video data, i.e., exceeding hundreds of hours, while ensuring that the computational overhead remains independent of the video length; 2) The index construction must operate in near-real-time, allowing the system to support timely event analytics; 3) The system should accommodate not only fact-based retrieval queries but also query-focused summarization and multi-hop queries, supporting open-ended analytics.
To this end, as depicted in Fig. 1, Avas system is composed of two primary components: near-real-time index construction and agentic retrieval and generation. Within each component, we introduce a set of techniques designed to effectively realize the established design principles.
In the index construction phase, our objective is to design an effective index while ensuring construction efficiency. To achieve this, we introduce the event knowledge graph (EKG) to structure video streams (§4.1). An EKG is a specialized form of a knowledge graph (KG) designed to represent and organize at the granularity of events and their interconnections. Recognizing that events unfold across varying temporal scales, we propose semantic chunking4.2) to extract meaningful events from video streams. Specifically, video streams are segmented into small, uniform chunks (e.g., 3-second intervals), and a small VLM, such as Qwen2.5-VL-7B, is periodically employed to generate detailed content descriptions for these chunks using carefully crafted prompts. Subsequently, neighboring chunks are merged into larger semantic chunks by identifying semantically equivalent textual descriptions with BertScore (Devlin et al., 2019). For each semantic chunk, the small VLM extracts entities and their relationships. Identical entities across different events are linked to ensure consistency and coherence. Ultimately, Avas facilitates the continuous construction of an EKG for a given video stream, regardless of its length, providing a comprehensive representation of semantic events, entities, and their interrelations in near-real-time on typical edge servers.
In the retrieval and generation process, we aim to leverage the constructed index for efficiently retrieving essential and minimal information, and to robustly generate the final answer based on the retrieved data. To achieve this, we first introduce the concept of tri-view retrieval. Specifically, a given query undergoes simultaneous retrieval across three dimensions: events, entities, and visual embeddings. This approach ensures the acquisition of comprehensive and relevant information pertaining to the query. To further support complex queries in L4 VA systems e.g., query-focused summarization and multi-hop queries, we propose an agentic searching mechanism. In particular, by utilizing the LLM as an agent, Avas proactively explores to retrieve additional information from events linked to those retrieved in earlier steps. Avas explores multiple pathways to gather information and formulates a response to the query based on the collected data. Finally, we introduce the thoughts-consistency strategy, which selects the most coherent and accurate final answer from multiple generated candidates.
4. Near-Real-Time Index Construction
4.1. Event Knowledge Graph
An Event Knowledge Graph (EKG) is a structured representation of events and their interconnections, linking entities, timestamps, locations, and other contextual information to offer a holistic understanding of events and their dependencies. By employing an EKG, the content of a video can be organized into a sequence of events, associating groups of entities with specific events and capturing their intricate relationships.
Although existing works, such as GraphRAG (Edge et al., 2024) and LightRAG (Guo et al., 2024), utilize knowledge graphs (KGs) to construct retrieval indices, we argue that EKGs are more suitable for video data. The rationale lies in the fundamental difference between the two: KGs focus on static entities (e.g., people, locations, concepts) and their attribute-based relationships, whereas EKGs prioritize modeling dynamic events and their spatiotemporal evolution.
Fig. 2 illustrates an example video alongside its corresponding KG and EKG. As depicted, the EKG effectively captures key events and their transitions, while representing entities with finer granularity within specific events. This enables EKG-based retrieval to support more sophisticated queries, such as event summaries, multi-hop temporal reasoning, and other complex analyses. In contrast, KGs, which only encapsulate entities across the entire video, lack the capability to fundamentally support such advanced queries.
Figure 2. An example of a constructed event knowledge graph and a knowledge graph from wildlife surveillance scenarios for video analytics.
Formally, we define our EKG $\mathcal{G}$ as follows:
(1) $\mathcal{G}=(\mathcal{E},\mathcal{U},\mathcal{R}),$
where =$\mathcal{E}=\{e_{i}\}_{i=1}^{|\mathcal{E}|}$ represents the temporally ordered set of events, $\mathcal{U}=\{u_{j}\}_{j=1}^{|\mathcal{U}|}$ denotes the entities extracted from the video within each event, and $\mathcal{R}=\mathcal{R}_{ee}\cup\mathcal{R}_{uu}\cup\mathcal{R}_{ue}$ encompasses three types of relationships: 1) temporal event-event relations $\mathcal{R}_{ee}$, such as before and after, which encode temporal logic constraints; 2) semantic entity-entity relations $\mathcal{R}_{uu}$, akin to the relationships found in conventional KGs; and 3) participation relations $\mathcal{R}_{ue}$, which associate entities with their contextual roles within specific events.
4.2. Semantic Chunking
Figure 3. Merging uniform chunks into semantic chunks guided by the pairwise BERTScore distribution.
To construct the EKG as an index, it is essential to extract events and their corresponding descriptions from videos. Although current VLMs demonstrate remarkable capabilities in event detection and transcription, their application in video analytics scenarios remains challenging.
On one hand, large VLMs, such as Qwen2.5-VL-72B, can achieve high accuracy in event detection and transcription, but their substantial computational overhead makes it difficult to process video streams in near-real-time, particularly on resource-constrained edge servers. On the other hand, small VLMs, such as Qwen2.5-VL-7B, offer reduced latency but suffer from performance degradation as the length of the video increases. Furthermore, both large and small VLMs are limited by their constrained context windows. To handle long video content, a common approach is to partition the content uniformly, a process known as chunking. However, events in videos naturally occur at varying and diverse temporal scales. Inaccurate chunking can disrupt the coherence of individual events, thereby increasing the difficulty for VLMs to accurately detect and transcribe them.
To address this, we propose a semantic chunking approach. The core idea involves processing a video stream in the following steps: First, we perform uniform buffering, e.g., dividing the video into fixed-length chunks of 3 seconds each. Next, a small VLM ,e.g., Qwen2.5-VL-7B, is employed to extract representative event descriptions from these chunks with proper prompts. Based on the generated event descriptions, we utilize a text embedding model, such as BERT (Devlin et al., 2019), to measure the similarity between neighboring events. Adjacent events with high similarity are then merged into a single event. Ultimately, this process enables Avas to partition the entire video into semantically meaningful chunks while simultaneously extracting their corresponding descriptions.
Particularly, an input video $V$ is initially divided into uniform chunks $c_{i}$, and a small VLM is employed to generate textual descriptions $d_{i}$ for each chunk $c_{i}$. Subsequently, the semantic similarity between any two uniform chunks is measured by computing the pairwise BertScore (Zhang et al., 2020) for $(d_{i},d_{j})$. Higher similarity scores suggest that the same event may occur across these chunks, making them candidates for semantic merging. Specifically, we adopt two criteria to determine whether certain uniform chunks can be merged into a single semantic chunk: 1) Within a semantic chunk, the similarity between any two uniform chunks must exceed a predefined threshold (e.g., 0.65 in our implementation); 2) After merging, the similarity between the boundaries of adjacent semantic chunks must fall below a sufficiently low threshold. Figure 3 illustrates the semantic chunking process, where a video initially divided into 18 uniform chunks is successfully merged into 9 semantic chunks. Once merged, the small VLM is further utilized to summarize each semantic chunk.
It is important to highlight, although pairwise BertScore computations are performed multiple times, Avas efficiently schedules these computations in parallel, leveraging the hardware parallelism (§6). Consequently, the semantic chunking process does not become a bottleneck in the near-real-time index construction phase, as detailed in §7.
4.3. Entity Extraction and Linking
In addition to extracting event information from videos, Avas also identifies entities and their relationships, as illustrated in Fig. 1. Similar to the approach in (Edge et al., 2024), we utilize a small VLM to extract entities and their relationships from videos using carefully designed prompts for each event.
The identified entities, however, tend to be highly redundant across events within the EKG. Such redundancy not only increases storage requirements but also hampers retrieval efficiency. Thus, it is necessary to de-duplicate and link these entities. Existing works (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025), which primarily focus on text-only RAG problems, typically rely on exact string matching strategies for entity de-duplication. However, in the context of video analytics and EKG, entities are independently extracted from each event by the VLM, leading to potential inconsistencies in entity descriptions for the same concept across different events, e.g.,, ”raccoon” and ”procyon lotor”.
To address this, Avas employs a text embedding model e.g., JinaCLIP (Koukounas et al., 2024), to encode all extracted entities into vector representations. Using embedding similarities as a metric, we apply a standard K-means clustering algorithm to group entities. This approach ensures that semantically similar entities are de-duplicated and linked by forming unified clusters. Subsequently, the centroid of each cluster is used as the representative for the linked entities.
Ultimately, the constructed EKG is stored in a database comprising five tables: events, entities, event-to-event relationships, entity-to-entity relationships, and entity-to-event relationships. Additionally, the raw video frames are vectorized using JinaCLIP (Koukounas et al., 2024) and linked to their corresponding events, enabling comprehensive retrieval in the following phase.
5. Agentic Retrieval and Generation
In the agentic retrieval and generation stage, our primary objectives are to effectively retrieve relevant information using the constructed EKG and to generate robust, contextually accurate responses based on the retrieved data. To achieve this, we introduce an agentic searching mechanism that explores multiple retrieval pathways within the index to gather the necessary information. Additionally, we propose a thought consistency strategy, enabling Avas to generate and evaluate multiple responses from different retrieval pathways, ultimately selecting the most appropriate one.
5.1. Tri-View Retrieval
Figure 4. An illustration of tri-view retrieval and borda counting on the retrieved events.
To comprehensively retrieve relevant information from the index for a given query, Avas employs a three-view retrieval process: the first view targets events, enabling retrieval at the event level to provide information for event summary-related queries. Specifically, the query is encoded using the text encoder JinaCLIP (Koukounas et al., 2024) and matched against the events table in the constructed EKG. The second view focuses on entities, offering insights into basic facts or item-specific queries. For this, we leverage the entity centroids extracted and aggregated as detailed in §4.3 to facilitate retrieval. The third view utilizes vision embeddings of raw video frames as complementary information. The retrieved entities and raw frames are subsequently linked to their associated events through the constructed EKG.
It is important to note that the retrieved events should be ranked. Ranking is not only crucial for filtering noise from the retrieved results but also essential for enabling agentic searching, as detailed in §5.2. A straightforward ranking method, such as similarity-based ranking, cannot be directly applied to Avas due to the retrieved events originating from three distinct views. To integrate these results, we propose to use a weighted Borda counting approach.
Fig. 4 illustrates the process of using Borda counting to integrate and rank retrieved events from the three views in Avas. Specifically, we select the top $K$ events from each view and rank them based on their calculated similarities within that view. Subsequently, the similarities of these $K$ events are normalized to compute their Borda scores:
(2) $s_{m}(e_{j})=\frac{\text{sim}_{m}(e_{j})}{\sum_{e_{k}\in\mathcal{E}_{m}}\text{sim}_{m}(e_{k})},$
where $\mathcal{E}_{m}$ represents the set of events retrieved from view $m$. The final Borda score for each event $e_{j}$ is then obtained by summing its scores across all views:
(3) $s(e_{j})=\sum_{m}s_{m}(e_{j}),$
Finally, the aggregated Borda scores $s(e_{j})$ are used to rank all retrieved events.
Figure 5. An example of agentic tree search with four actions and a depth of three, yielding 13 distinct pathways for information gathering and response generation.
5.2. Agentic Searching on Graph
The retrieved events mentioned above can be directly utilized to generate the final answer. However, to support complex queries, such as query-focused summaries and multi-hop queries, Avas searches for additional relevant information by leveraging the relationships between events and entities within the constructed EKG. To enable efficient exploration, we propose the agentic searching on graph approach.
This approach is inspired by human behavior when retrieving and reasoning about information in videos. Specifically, individuals typically begin by identifying key clips based on retrieval keywords, then gather additional context by exploring clips preceding or following the identified ones. To achieve a more comprehensive understanding, they may also refine their search by using alternative keywords, iterating through this process as needed. Similarly, in our agentic searching process, we define the agentic action space as follows:
Forward (F): this action extends the current retrieval by including temporally subsequent events on the EKG for all events in the event list. It reflects the natural tendency of humans to seek forward narrative progression when trying to understand what happens next or how a situation evolves over time.
Backward (B): complementing the forward action, this action retrieves temporally preceding events, enabling a backward exploration of the narrative to uncover prior context or causal factors.
Re-query (RQ): this action generates a new query represented by a list of keywords via a LLM and retrieves complementary events as outlined in §5.1. It reflects the human tendency to gather information from multiple perspectives to achieve a more comprehensive understanding.
Summary and answer (SA): this action utilizes the descriptions of the retrieved events from the EKG and generates the response to the specific query by employing a LLM.
By leveraging these predefined agentic actions, we design the agentic searching process as a tree search framework. The search begins with an initial retrieval based on the original query, producing a list of relevant events that serves as the root node of the search tree. At each depth level, a one-step rollout of all four predefined actions is performed for each active node. If the SA action is reached, the corresponding search path is terminated. The rollout process continues until the maximum tree depth is reached. This tree search process explores diverse pathways to gather information from the EKG; As iilustrated in Fig. 5, a tree with a depth of 3 would result in 13 distinct information-gathering paths and corresponding answers. All generated answers are evaluated, and the optimal one is selected using the thought consistency method, which is detailed in §5.3.
A practical issue in the tree search process is the exponential growth in the number of retrieved events as the tree depth increases. This not only introduces computational overhead but also results in the accumulation of noisy or irrelevant information. To mitigate this, we use a length constraint on the maintained event list during the search process, i.e., 16 in our implementation. When the number of retrieved events exceeds this limit, we employ a drop strategy to discard less relevant events based on their rankings described in §5.1.
5.3. Consistency Enhanced Generation
During the agentic tree search, multiple candidate answers are generated at SA nodes across different pathways. To determine the final answer, it is necessary to either select or synthesize from these candidates. A straightforward approach would be majority voting. However, due to the diversity of retrieval paths, only a small subset of these nodes is likely to access the essential information with minimal noise, producing high-quality answers. To this end, we introduce the thoughts-consistency mechanism to identify and select the most reliable final answer.
At each SA node, instead of generating the answer a single time, we repeatedly generate answers multiple times using a Chain-of-Thought (CoT) prompting scheme. Following the principle of self-consistency (Wang et al., 2023), correct answers are more likely to emerge consistently across multiple valid reasoning trajectories during repeated generations. Specifically, we evaluate the consistency not only across the generated answers but also within their associated CoT traces. Nodes demonstrating strong internal coherence, where the reasoning process aligns logically with the conclusion, are assigned higher scores.
To formalize this process, we propose a scoring framework that integrates both answer agreement and thought consistency. At each SA node, we perform $n$ rounds of sampling using a temperature setting between $0.5$ and $0.7$, resulting in a set of $n$ candidate outputs denoted as $\{(a_{i},r_{i})\}_{i=1}^{n}$, where $a_{i}$ is the answer and $r_{i}$ is the associated reasoning trace.
Let $\mathcal{A}=\{a^{(1)},a^{(2)},\dots,a^{(T)}\}$ be the set of unique answers among the $n$ samples, where $T$ is the number of distinct answers. The answer agreement score $S_{a}^{(t)}$ for a candidate answer $a^{(t)}$ is defined as the proportion of times it appears in the samples:
(4) $S_{a}^{(t)}=\frac{|\{i\mid a_{i}=a^{(t)}\}|}{n}$
The thought consistency score $S_{r}^{(t)}$ for $a^{(t)}$ is computed as the average BERTScore between all pairs of reasoning traces associated with $a^{(t)}$:
(5) $S_{r}^{(t)}=\frac{2}{k(k-1)}\sum\limits_{1\leq i<j\leq k}\text{BERTScore}(r_{i},r_{j}),$
where $k$ is the number of times $a^{(t)}$ appears in the $n$ samples. The final score for each candidate answer combines these two components:
(6) $S_{\text{final}}^{(t)}=\lambda S_{a}^{(t)}+(1-\lambda)S_{r}^{(t)},$
where $\lambda\in[0,1]$ is a weighting parameter controlling the trade-off between answer agreement and thought consistency. In our implementation, we set it to 0.3, the parameter tuning would be discussed in §7.4.3.
For each SA node, the candidate answer with the highest $S_{\text{final}}^{(t)}$ is selected as its definitive response. To enhance the reliability of this final answer, we propose an additional agentic action, Check Frames and Answer (CA). This action retrieves the raw video frames associated with the events from the EKG and utilizes the VLM to generate a refined response to the specific query. By doing so, this action effectively supplements any missing information relevant to the query that may have been overlooked during the construction phase.
Specifically, after ranking all candidate answers from the SA nodes using the consistency-enhanced scoring mechanism, the top-2 nodes with differing answers are selected. The video frames corresponding to their retrieved events are extracted, and the VLM is prompted to generate a new response by directly attending to the visual evidence. Furthermore, the thought-consistency mechanism is applied to the CA nodes to bolster the reliability of the final generated answer.
6. Implementation
Avas uses Qwen2.5-VL-7B for constructing EKGs, Qwen2.5-32B for SA, and Gemini-1.5-Pro for CA. Local VLMs and LLMs are deployed using LMDeploy(Contributors, 2023) to enable accelerated inference. Additionally, Avas adopts batch inference for several key stages—including description generation, description merging, entity extraction, and tree search—to improve efficiency and maximize GPU utilization. For text and vision embedding, we utilize JinaCLIP(Koukounas et al., 2024). And we employ BERTScore with the deberta-xlarge-mnli(He et al., 2020) checkpoint. The storage of EKG and vector representations is based on the implementation of (Guo et al., 2024), upon which we make further modifications to suit the specific requirements of Avas.
During the EKGs construction stage, we carefully design prompts to guide the extraction of structured information from video. For general-purpose video understanding, we employ a unified prompt that avoids introducing bias or prior assumptions:
”…Your task is to extract and provide a detailed description of the video segment, focusing on all key visible details…”.
For scenario-specific videos, we design the prompts to emphasize scenario-relevant information. For example, in the case of wildlife surveillance scenario, key information may includes the timestamps of the recording, animal activities (e.g., presence, species, number, specific behaviors, etc.), and environmental changes.
7. Evaluation
7.1. Evaluation Settings
Figure 6. The achieved accuracy of Avas and various baselines on the LVBench, VideoMME-Long, and Avas-100 benchmarks.
7.1.1. Benchmark
Avas is evaluated on two widely used public long-video benchmark and one ultra-long video benchmarks proposed by us, covering a broad range of video scenarios and problem types.
LVBench(Wang et al., 2024b) stands out among publicly available benchmarks for its exceptionally long average video duration, approximately 4100 seconds per video. It comprises 103 videos with a total of 1549 questions, covering six distinct video domains and addressing six task types including temporal grounding, summarization, and reasoning.
VideoMME-Long(Fu et al., 2024) is a subset of the VideoMME benchmark, focusing on videos exceeding 20 minutes in duration, with an average length of 2400 seconds. Comprising a total of 300 videos and 900 questions, the benchmark covers a wide range of video themes, spanning 6 primary visual domains with 30 subfields to ensure broad scenario generalizability, and includes 12 distinct task types.
Avas-100 is proposed by us, which is an ultra-long video benchmark specially designed to evaluate video analysis capabilities Avas-100 consists of 8 videos, each exceeding 10 hours in length, and includes a total of 120 manually annotated questions. The benchmark covers four typical video analytics scenarios: human daily activities, city walking, wildlife surveillance, and traffic monitoring, each scenario contains two videos. The human daily activity scenario features egocentric videos selected and stitched from the Ego4D(Grauman et al., 2022). City walking and wildlife surveillance videos are curated from publicly available recordings on YouTube, capturing urban exploration and animal monitoring respectively. Traffic monitor videos are composed from clips in the Bellevue Traffic Video Dataset (Bhardwaj et al., 2022). All questions are carefully designed by human annotators, who also provide reference answers as the ground truth. In addition, GPT-4o is utilized to generate plausible distractor options. The accuracy is evaluated by analyzing Avas’s responses to multiple-choice questions included in the benchmarks.
7.2. Baselines
We conduct a comprehensive comparison between Avas and a wide range of baseline models, encompassing both mainstream VLMs and specialized Video-RAG methods. The VLM baselines include GPT-4o (Achiam et al., 2023), Gemini-1.5-Pro (Team et al., 2023), Phi-4-Multimodal (Abdin et al., 2024b), Qwen2.5-VL-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024), and LLaVA-Video-7B (Zhang et al., 2024a). Each of these models is evaluated with two typical strategies: uniform sampling and vectorized retrieval, where a CLIP-based retriever selects the top-K relevant frames based on the user query. In addition to VLMs, we benchmark Avas against SOTA Video-RAG frameworks, including VideoTree(Wang et al., 2024c), VideoAgent(Wang et al., 2024d), DrVideo(Ma et al., 2024), and VCA(Yang et al., 2024). Among these, VideoTree, VideoAgent, and VCA are built upon GPT-4o, while DrVideo leverages GPT-4.
7.3. Overall Evaluation
7.3.1. Overall Performance
Fig. 6 illustrates the overall accuracy achieved by Avas compared to various baselines on the LVBench, VideoMME-Long, and Avas-100 benchmarks. Across all three benchmarks, Avas consistently outperforms the baselines. Specifically, on LVBench, Avas delivers a remarkable 16.9% improvement, while on VideoMME-Long, it advances the SOTA by approximately 5.2%. On the Avas-100 benchmark, Avas achieves an accuracy of 75.8%, significantly surpassing all competing methods.
In detail, compared to video-RAG methods, Avas achieves improvements of 21% and 7.8% on LVBench and VideoMME-Long, respectively. When compared to vectorized retrieval-based methods, Avas demonstrates gains of 16.9% on LVBench and 20.8% on Avas-100. Furthermore, against uniform sampling baselines, Avas improves performance by approximately 19.6% and 26.9% on LVBench and Avas-100, respectively.
Notably, on Avas-100, when evaluated with extremely long videos, Avas maintains robust performance, whereas the baselines degrade significantly. This highlights the effectiveness of Avas in handling L4 video analytics tasks.
Figure 7. The accuracy achieved by Avas and the baselines across typical query categories on LVBench: Temporal Grounding (TG), Summarization (SU), Reasoning (RE), Entity Recognition (ER), Event Understanding (EU), and Key Information Retrieval (KIR).
7.3.2. Performance on Different Query Categories
We also evaluate the accuracy achieved by Avas across typical query categories on LVBench. As illustrated in Fig. 7, our approach achieves improvements of 16%, 5.3%, 35.6%, 21.2%, 17.5%, and 18.9% across six key task types: Temporal Grounding, Summarization, Reasoning, Entity Recognition, Event Understanding, and Key Information Retrieval, respectively, compared to the uniform sampling and vectorized retrieval baselines powered by Gemini-1.5-Pro. Notably, Avas demonstrates particularly strong performance on reasoning tasks, which require identifying causal relationships between events and linking preceding and succeeding events within the video. This highlights Avas’s ability to effectively locate and extract critical information from long videos, thereby enabling advanced L4 video analytics systems.
Figure 8. The accuracy achieved by Avas and baselines across three benchmarks when utilizing different LLMs and VLMs.
7.3.3. Performance under Different Configurations
Fig. 8 shows the performance of Avas using different models configurations for SA and CA. For SA, two models were used: Qwen2.5 14B and 32B. For CA, two models were used: Qwen2.5-VL-7B and Gemini-1.5-Pro. The results show that across the three benchmarks, Avas using Gemini-1.5-Pro for CA achieved improvements of 18.9%, 5.2%, and 20.8% respectively compared to the best baseline result using the same model, while using Qwen2.5-VL-7B yielded improvements of 13%, 7.2%, and 15% respectively, fully demonstrating the effectiveness of our method. Notably, even when only using Qwen2.5-32B and Qwen2.5-7B based on the textual content from EKG without accessing raw frames, Avas can surpass the performance of Qwen2.5-VL-7B on the three benchmarks and also outperform most models shown in Figs. 6(a), 6(b), and 6(c).
7.3.4. Performance on Different Video Lengths
Figure 9. The accuracy achieved by Avas and the baselines across varying video lengths via concatenating videos from LVBench.
To evaluate the robustness of Avas with respect to video length, we conducted experiments on videos of varying durations. Specifically, sequences of 3.3, 6.6, and 10 hours were created by concatenating videos from the VideoMME-Long benchmark. Performance was measured using identical questions across these varying video lengths. As illustrated in Fig. 9, both Qwen2.5-VL-7B and Gemini-1.5-Pro baselines exhibit significant performance degradation as video length increases. When extended to 10 hours, their performance declines by 4.6% and 8.2%, respectively, under the uniform sampling method, compared to the original VideoMME-Long benchmark. For the vectorized retrieval setting, the performance drops are 4.6% and 5.5%, respectively. These results highlight the limitations of these methods in scaling effectively with increasing video length. In contrast, Avas consistently maintains stable performance across all video lengths, underscoring its robustness and scalability in handling video data of any duration.
7.3.5. System Overhead
Avas is designed to enable the near real-time construction of EKGs. As shown in Fig. 10, we measured the average processing speed (in FPS) of Avas while constructing EKGs from LVBench videos across various hardware platforms, with the input video stream fixed at 2 FPS. On 2 $\times$ A100 GPUs, Avas achieved an impressive processing speed of 6.7 FPS, significantly exceeding the input stream rate. On a single RTX 4090, a typical edge server hardware, Avas maintained a processing speed of 4.4 FPS, still well above the input frame rate. Even on a single RTX 3090, Avas performed effectively, achieving 2.5 FPS. This performance demonstrates its capability to support efficient, near real-time EKG construction for L4 video analytics. The overhead during the retrieval and generation phases will be discussed further in §7.4.
To evaluate the generation overhead of Avas, we performed a latency breakdown of the key stages in the generation pipeline, as detailed in Tab. 2. The results indicate that a single retrieval operation is very swift, requiring only 0.44 seconds. This efficiency stems from the fact that this step involves vector matching between information established during the graph construction phase. The tree search stage accounts for a substantial portion of the overall generation latency. When employing Qwen2.5-14B, this stage exhibits a latency of 101.5 seconds. This duration increases correspondingly to 174.2 seconds when switching to the larger Qwen2.5-32B, which has a greater number of parameters. For check and answer (CA) stage, Qwen2.5-VL-7B demostrates a latency of 85.2 seconds, while the API-based Gemini-1.5-Pro shows a latency of 50.2 seconds.
Figure 10. Total index construction overhead evaluated on various types of typical edge server hardware.
Stage Model Latency(s)
Retrieval JinaCLIP 0.44
Tree Search Qwen2.5-14B 101.5
Qwen2.5-32B 174.2
Check and Answer Qwen2.5-VL-7B 85.2
Gemini-1.5-Pro 50.2
Table 2. The latency breakdown for the retrieval and generation phase of Avas.
7.4. Ablation Evaluation
We randomly sampled 20 videos and 305 corresponding questions from LVBench for our ablation study. All ablation experiments were conducted on 2 $\times$ A100 GPUs.
7.4.1. Different Index Construction Methods
Method Acc. Overhead(h)
MiniRAG 28.1 3.49
LightRAG 30.6 3.52
Avas 39.7 0.31
Table 3. The achieved accuracy and construction overhead evaluated when using EKG and KG as index in Avas and baseline models on the subset of LVBench. The total video duration is around 1.2 hours.
We compare Avas’s Event Knowledge Graph (EKG) construction method with two representative knowledge graph-based construction methods: LightRAG (Guo et al., 2024) and MiniRAG (Fan et al., 2025). Since both of them only support text-only construction, we use the full set of descriptions obtained through the semantic chunking (§4.2) as their input textual corpus. We use Qwen2.5 7B to construct EKG and KG for Avas and baselines, respectively. In the retrieval and generation phase, we use the same LLM, Qwen2.5 14B and the same settings, e.g., maximum number of tokens of retrieved events or entities.
As shown in Table 3, Avas demonstrates a significant performance advantage over the baselines. Specifically, it achieves 11.6% higher accuracy than MiniRAG and 9.1% higher accuracy than LightRAG. Crucially, this improved performance comes with substantially less construction overhead, requiring only 0.31 hours compared to 3.49 and 3.52 hours for the baselines. The rationale is that baselines construct KG based on massive uniform chunks, while Avas utilizes the semantic chunks. This substantial gap in both effectiveness and efficiency highlights that Avas’s EKG construction method not only yields higher-quality knowledge representations but also drastically reduces the time needed to build the graph.
7.4.2. Different Tree Search Depths
Method Tree Search Depth
1 2 3 4
Avas(Qwen2.5 14B) 34.1 36.1 40.9 39.5
Avas(Qwen2.5 14B + Qwen2.5VL 7B) 49.3 52.1 53.8 50.2
Avas(Qwen2.5 14B + Gemini-1.5-Pro) 54.2 58.4 61.5 52.7
Tree Search Overhead(s) 6.7 27.3 90.1 370.3
Table 4. The achieved accuracy and overhead when applying different tree search depths in the agentic search of Avas evalauted on the subset of LVBench.
We also evaluate the achieved accuracy and overhead applying different settings in the retrieval phase, i.e., tree search depth. The effect of tree depth lies in a crucial trade-off: while shallower depths may struggle to retrieve comprehensive information, increasing the depth allows access to richer information from deeper nodes. However, this comes with a significant increase in tree search overhead, and the information from deeper levels can introduce more noise, potentially negatively impacting the final generation quality. Table 4 presents the results comparing different tree search depths on performance and tree search overhead. As shown, performance generally increases with increasing tree depth up to a certain point. Specifically, for all three Avas configurations evaluated, the highest accuracy is achieved at a tree search depth of 3. Accuracy decreases when the depth is further increased to 4, suggesting that excessive depth leads to the retrieval of detrimental noise or irrelevant information, outweighing the benefit of additional context. Conversely, the tree search overhead increases sharply with depth. Expanding the search from depth 1 (6.7s) to depth 2 (27.3s) incurs a moderate increase. Comparing the accuracy improvements and the overhead increase, a tree search depth of 3 offers the optimal balance.
7.4.3. Different Consistency Evaluation Settings
For consistency-enhanced generation, Avas incorporates two key parameters: $\alpha$, which governs the balance between the contributions of thought consistency and answer consistency, and the number of generations for self-consistency evaluation. Fig. 11(a) illustrates the impact of varying $\alpha$ values on the accuracy achieved by Avas. Notably, the optimal performance is observed when $\alpha$ is set to $0.3$, highlighting the importance of jointly considering both intermediate thought consistency and final answer consistency to ensure robust results. As depicted in Fig. 11(b), the accuracy of Avas gradually improves as the number of self-consistency iterations increases. However, this improvement comes at the expense of significantly higher computational overhead. For example, increasing the self-consistency iterations from 8 to 16 yields only a 0.9% accuracy gain, while nearly doubling the computational cost. This demonstrates a clear trade-off between marginal accuracy improvements and resource efficiency. Balancing this trade-off, we adopt 8 self-consistency iterations in the implementation of Avas, ensuring a practical balance between performance and computational overhead.
Figure 11. The performance of Avas under varying consistency evaluation settings on the subset of LVBench.
8. Limitations and Future Work
There are also limitations in the current design of Avas, which we explore for future work. Specifically: 1) The existing agentic retrieval and generation mechanism relies on a fixed tree-search strategy based on the Monte Carlo approach. While effective, this method is computationally expensive. The trajectories collected during the search process could be leveraged as training data to develop a model capable of dynamically selecting optimal search actions and depths based on the query and context. 2) Although the integrated VLM demonstrates robust general video understanding and reasoning capabilities, it may encounter challenges in certain specialized visual tasks, such as precise object counting. Incorporating lightweight, task-specific vision models as tools within the system could improve accuracy for such queries. Our future work will focus on enabling the VLM, functioning as an autonomous agent, to intelligently invoke these specialized tools, thereby addressing its limitations in handling specific tasks.
9. Conclusion
This paper presents Avas, an advanced L4 video analytics system powered by VLMs. Avas enables comprehensive understanding and open-ended query analysis of large-scale, long-duration video data, overcoming the constraints of existing video analytics systems that are predominantly tailored to specific, pre-defined tasks. The system introduces novel designs, including near-real-time Event Knowledge Graph (EKG) index construction and an agentic retrieval and generation mechanism, facilitating efficient organization and analysis of extended video content to address complex queries. We demonstrate Avas’s superior performance on public video understanding benchmarks, as well as on our newly proposed benchmark, Avas-100, specifically designed to evaluate video analytics tasks.